The probability of making the basketball team is 28%, the probability of making the softball team is 34%, and the probability of making one or the other is 50%. What is the probability of making both teams?
Answer by stanbon(75887)

The probability of making the basketball team is 28%, the probability of making the softball team is 34%, and the probability of making one or the other is 50%. What is the probability of making both teams?
------------------
P(B and S) = P(B)+P(S)-P(B or S)
------
P(B and S) = 0.28+0.34 - 0.50 = 0.12
----------------
Cheers,
Stan H.
-----------------